Ok diff (a library for differential operators by Okutani) Manual

Edition : auto generated by oxgentexi on 23 February 2009

OpenXM.org

1 Differential equations (library by Okutani)

The following functions will be imported. odiff act, odiff act appell4, odiff op appell4, odiff op fromasir, odiff op toasir, odiff op tosm1, odiff poly solve, odiff poly solve appell4, odiff poly solve hg1, odiff rat solve.

English manual has not been written.

Index 2

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 Differential equations (library by Okutani) . . . . . . . . . . . . . . 1 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

1 Differential equations (library by Okutani) . . 1 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2